# Non-transitive pseudo-Anosov flows

### v1

@@ Abstract.
We study (topological) pseudo-Anosov flows from the perspective of the associated group actions on their orbit spaces and boundary at infinity. We extend the definition of Anosov-like action from [BFM22] from the transitive to the general non-transitive context and show that one can recover the basic sets of a flow, the Smale order on basic sets, and their essential features, from such general group actions. Using these tools, we prove that a pseudo-Anosov flow in a $3$ manifold is entirely determined by the associated action of the fundamental group on the boundary at infinity of its orbit space. We also give a proof that any topological pseudo-Anosov flow on an atoroidal 3-manifold is necessarily transitive, and prove that density of periodic orbits implies transitivity, in the topological rather than smooth case.
@@ 1. Introduction
This paper is motivated by the structure theory and classification of pseudo-Anosov flows in dimension 3, including in the topological setting where flows may not have strong stable and unstable manifolds. Such flows are important examples in hyperbolic dynamics and low-dimensional topology, connected to many geometric-topological structures on 3-manifolds, and the problem to classify such flows dates back to Smale.
A pseudo-Anosov flow $\phi$ on a 3-manifold $M$ gives rise to an orbit space $\mathcal{O}_{\phi}$, a topological plane with two transverse, possibly singular foliations which admits a natural action of $\pi_{1}(M)$ by homeomorphisms. Following work of Barbot [Bar95a], this group action up to conjugacy uniquely determines the flow up to orbit equivalence. One result of this paper is a further reduction of dimension:
Theorem 1.1 (Circle actions classify flows).
The action of $\pi_{1}(M)$ on $S^{1}$, arising from the compactification of the orbit space by a circle at infinity, completely determines a pseudo-Anosov flow up to orbit equivalence.
In fact we prove a more general result – see Theorem 1.10.
More broadly, Barbot’s theorem indicates that all features of a pseudo-Anosov flow and its orbit equivalence class should, in theory, correspond to features of the orbit space with the action of $\pi_{1}(M)$. Building such a “dictionary” between flows and their orbit spaces has recently become an essential tool in the subject. Early work of Barbot and Fenley used this perspective towards classifying pseudo-Anosov flows on graph manifolds [Bar96, BF13, BF15], among other results. Recently, the orbit space perspective was used in [BFM22, BFM23] to obtain a complete algebraic invariant of transitive pseudo-Anosov flows (on any given manifold $M$) via the dynamics of $\pi_{1}(M)$ on $\mathcal{O}_{\phi}$. More generally, that work studied transitive Anosov-like actions on bifoliated planes, of which the actions on orbit spaces are examples, giving a complete invariant to classify such actions up to conjugacy.
@@
This paper develops the theory of Anosov-like actions on bifoliated planes, with no reference to transitivity, thus building a toolkit to study general pseudo-Anosov flows from their orbit spaces. For instance, we give a proof, purely from the orbit space perspective, that any (topological) pseudo-Anosov flow or topological Anosov flow on an atoroidal 3-manifold is necessarily transitive. This fills an apparent gap in the literature, since the existing proofs for smooth Anosov or pseudo-Anosov flows, due to Brunella [Bru93] and Mosher [Mos92], use Smale’s spectral decomposition theorem, which is only fully proved in the smooth setting. We also show explicitly that many essential, classical dynamical-topological features, such as basic sets and separating tori, can be detected concretely from the orbit space alone; building the dictionary correspondence between flows and orbit spaces. We also prove global structural results and characterize transitive versus nontransitive flows. This resolves several questions left open in [BFM22].
To precisely state our results, we first give the definition of an Anosov-like action.
@@
A bifoliated plane is a topological plane $P$ with two transverse topological foliations $\mathcal{F}^{\pm}$, possibly with isolated prong singularities.
@@
For the results of this article, Axiom (A5) could be weakened slightly, requiring that we only forbid such quadrilaterals that do not contain any fixed points of elements of $G$ (see Remark 4.19).
Remark 1.3.
As a consequence of the axioms, if a leaf $l$ has nontrivial stabilizer, then all nontrivial elements of the stabilizer have a common, unique fixed point in $l$. For singular leaves, this is a consequence of Axioms (A1) and (A3). In the nonsingular case, this follows from Axiom (A1) together with Solodov’s Theorem (see [BFM22, Lemma 2.6]).
Proposition 2.2 below shows that the orbit space action from a pseudo-Anosov flow satisfies these axioms; however these are not the only examples.
These axioms have surprisingly strong consequences, both for the topological structure of the bifoliated plane, and the global and local dynamics of the action of $G$. For example, we show (following Barbot) that a bifoliated plane admitting an Anosov-like action either has prongs or nonseparated leaves, is trivial, or has a special form called skew (Proposition 2.8). There is also a relationship between dynamics and group structure:
Theorem 1.4.
@@
The proof does not require any 3-manifold topology (beyond showing that the orbit space of a flow is Hausdorff), nor require any smoothness of the flow, so applies also to topological (pseudo)-Anosov flows in the sense of [BFP23, Def. 2.1] or [AT24, Def. 5.9]. We also show
@@
Again, this is well known for smooth Anosov flows, but to our knowledge not proved in the topological and pseudo-Anosov setting.
For these latter results, we need to develop the theory of Smale classes.
Smale classes and chains. A Smale class for an Anosov-like action is a $G$-invariant set generalizing the notion of a basic set for an Anosov flow. These inherit a natural partial order, which we denote by $\leq_{G}$ (see Definition 3.3). We show
@@
This answers a problem from [BFM22, Remark 2.4], see Corollary 1.11.
By [Bru93], when $M$ is orientable and compact, any two basic sets of an Anosov flow are separated by embedded tori in $M$ transverse to the flow. We describe the counterpart structure in the orbit space, that we call Smale chains and show
@@
See Theorem 4.23 for a more precise statement, in which we clarify what is meant by separating.
Together with Theorem 5.5, this shows that, provided an Anosov-like action is cocompact, distinct Smale classes can be separated by $\mathbb{Z}^{2}$-invariant wandering chains.
Our framework also applies to pseudo-Anosov flows, where the situation is more subtle. For pseudo-Anosov flows, wandering chains separate distinct non-wandering components, which can be strictly finer than chain-recurrent components (see Proposition 8.8) which are the components separated by Lyapunov functions.
The boundary at infinity. A bifoliated plane comes equipped with a natural compactification by a boundary at infinity (see [BFM22, Bon24, Fen12, Mat82]) to which any foliation-preserving action by homeomorphisms extends. This compactification has become an important tool for the study of group actions on such planes and, as a special case, the study of pseudo-Anosov flows. It was shown in [BFM22] that for transitive Anosov-like actions, the induced action at infinity uniquely determines the action on the interior. Here, using tools from [BBM24], we show this is true generally, with an independent proof.
@@
Let $G$ be an Anosov-like action on a bifoliated plane $P$. The induced action on $\partial P$ uniquely determines the action of $G$ up to conjugacy.
Consequently, pseudo-Anosov flows on 3-manifolds are determined up to orbit equivalence by the action of $\pi_{1}(M)$ on $\partial P$ as stated in Theorem 1.1.
The proof of this theorem has two ingredients: One is the work of the same authors done in [BBM24], which allows to uniquely recover the bifoliated plane (and, by uniqueness, also the group action) from relatively “sparse” data on the boundary (Corollary E of [BBM24], restated as Theorem 7.5 below). The second ingredient, obtained in this article, consists in proving that when the foliations are not $\mathbb{R}$-covered, the non-corner fixed leaves are dense in $P$. The endpoints of such leaves are the sparse – it is a countable set – but sufficiently rich so that we can apply the result of [BBM24]. A non-corner fixed leaf is a leaf $l$ of $\mathcal{F}^{\pm}$, fixed by some nontrivial element $g\in G$, and such that the fixed point $x$ of $g$ on $l$ is not the corner of any lozenge (see Definition 2.5). Such non-corner fixed points and fixed leaves can be “seen” by the action of $G$ on the circle at infinity: an element $g$ fixes a non-corner leaf if and only if the subgroup generated by $g$ has exactly $4$ fixed points on $\partial P$ such that two are attracting and two are repelling (see Propositions 7.1 and 7.2). Thus, having two actions that are conjugated on the circle at infinity forces the non-corner fixed leaves for both actions to be “the same”, and the result of [BBM24] then gives the conjugacy.
Thus, the technical work toward the proof of Theorem 1.10 is to show the density of non-corner fixed leaves. This is obtained in Corollary 6.8. In the case of topologically transitive Anosov-like actions, this is comparatively easy, and was obtained in [BFM22]. In the non-transitive case however, it relies on the description of Smale classes and Smale chains that we obtain here.
@@
The axioms for transitive Anosov-like actions. Theorem 1.8 answers the question posed in [BFM22, Remark 2.4], and in fact shows the axioms given there for transitive Anosov-like actions are redundant and can be simplified. To explain this, we recall that [BFM22] gave a list of strictly stronger axioms than those in Definition 1.2 for groups acting on bifoliated planes, capturing the behavior of transitive pseudo-Anosov flows. In the transitive case, Axiom (A2) does not appear, and in its place we have the following two axioms.
@@
We point out that we also weakened slightly Axiom (A4) compared to [BFM22]. We show in Proposition 2.10 that the weaker axiom of this article imply the stronger one of [BFM22].
@@
Section 2 contains background on bifoliated planes and Anosov-like actions. We extend some results from [BFM22] to the nontransitive case, as well as introduce definitions and a toolkit used later in this work.
@@
Section 5 gives the proof of Theorems 1.4 and 1.5 and improves Theorem 1.9 in the case of cocompact actions.
@@ 2. Preliminaries: Anosov-like actions and bifoliated planes
Our definition of Anosov-like actions on bifoliated planes is motivated by the dynamics of the action of the fundamental group of a 3-manifold on the orbit space of a pseudo-Anosov flow. If $\phi$ is a flow on a manifold $M$, the orbit space $\mathcal{O}_{\phi}$ of $\phi$ is the quotient of $\widetilde{M}$ by orbits of the lift of $\phi$ to $\widetilde{M}$. Fenley [Fen94] and Barbot [Bar95a] showed that, if $\phi$ is a (pseudo)-Anosov flow [^1] The case of pseudo-Anosov flow was treated by Fenley–Mosher in [FM01]. on a 3-manifold, $\mathcal{O}_{\phi}$ is a topological plane. The weak stable and unstable foliations descend to transverse 1-dimensional singular foliations ±$\mathcal{F}^{\pm}(\phi)$ on $\mathcal{O}_{\phi}$ preserved by the action of $\pi_{1}(M)$.
@@
There are two common notions of Anosov and pseudo-Anosov flows in the literature: smooth and topological (pseudo)-Anosov flows (see, e.g., [AT24]). These notions are only known to coincide, up to orbit equivalence, for transitive pseudo-Anosov flows ([Sha21, AT24]). The following proposition applies even for the more general topological version of the definition.
@@ 2.1. Structure of the bifoliated planes
Following work of Barbot and Fenley, there is a well developed structure theory for the orbit spaces of Anosov flows. Several of these foundational results were established for bifoliated planes admitting (transitive) Anosov-like actions in [BFM22]. Below, we recall some of this structure theory and establish necessary extensions to our setting.
@@
A ray of ±$\mathcal{F}^{\pm}(x)$ is a properly embedded copy of $[0,\infty)$ in ±$\mathcal{F}^{\pm}(x)$ based at $x$ [^2] Elsewhere these are also called half-leaves, we prefer the term ray because it is more accurate in the case of singular points.
@@
In other words, a closed lozenge is homeomorphic to a rectangle with two opposite corners removed, with the segments of $\mathcal{F}^{+}$-leaves (resp. $\mathcal{F}^{-}$) sent to horizontal (resp. vertical) segments.
A point that is not the corner of any lozenge is called a noncorner point. If this point is fixed by some element $g\in G$, we call this a non-corner fixed point. These will play an important role later on. Lozenges may share corners, leading to what is called a chain of lozenges.
@@
A chain of lozenges is a union of closed lozenges that satisfies the following connectedness property: for any two lozenges $L,L^{\prime}$ in the chain, there exist lozenges $L_{0},\dots,L_{n}$ in the chain such that $L=L_{0}$, $L^{\prime}=L_{n}$, and, for all $i$, $L_{i}$ and $L_{i+1}$ share a corner (and may or may not share a side).
@@
The following global structure result is proved for transitive Anosov-like actions in [BFM22, Theorem 2.16], following the outline of [Bar95a, Théorème 4.1]. It also holds for general Anosov-like actions, with minor modifications to the proof, which we include below.
@@
    The bifoliated plane $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ is trivial.
@@
    The bifoliated plane $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ is homeomorphic to the diagonal strip bounded by $y=x-1$ and $y=x+1$ in $\mathbb{R}^{2}$ with the horizontal and vertical foliations. Such planes are called skewed.
@@
We describe the outline of the proof, for completeness, and describe the required modification from the proof given in [BFM22, Theorem 2.16].
@@
Let $\mathcal{L}^{-}$ and $\mathcal{L}^{+}$ denote the leaf spaces of $\mathcal{F}^{-}$, $\mathcal{F}^{+}$ respectively. Suppose that $P$ has no singular points and $\mathcal{L}^{+}$ is Hausdorff, since it is a simply connected 1-manifold it is thus homeomorphic to $\mathbb{R}$. We need to show in this case that $\mathcal{L}^{-}\cong\mathbb{R}$ and the plane is either skewed or trivial. Following the argument from [BFM22, Theorem 2.16] verbatim, one shows that, either both leaf spaces are $\mathbb{R}$ and the plane is trivial, or each leaf $l\in\mathcal{L}^{-}$ meets a bounded subset of leaves of $\mathcal{L}^{+}$. Fix an identification of $\mathcal{L}^{+}$ with $\mathbb{R}$ and for $l\in L^{-}$, let $\alpha(l)$ and $\omega(l)$ denote the infimum, and supremum, respectively, of the sets of leaves of $\mathcal{F}^{+}$ intersected by $l$.
We claim any leaf $l$ of $\mathcal{F}^{-}$ separates $\alpha(l)$ and $\omega(l)$ in $P$. This claim implies that $\mathcal{L}^{-}$ is Hausdorff, hence homeomorphic to $\mathbb{R}$, and thus $P$ can be seen as a subset of the plane obtained as the product of the leaf spaces +×$\mathcal{L}^{+}\times\mathcal{L}^{-}\simeq\mathbb{R}^{2}$ bounded by the graphs of the function $\alpha$ and $\omega$, as in [Fen94, Theorem 3.4], this is easily seen to be homeomorphic to the skewed plane.

### v2

# Nontransitive pseudo-Anosov flows

@@ Abstract.
This paper describes a structure theory for (topological) pseudo-Anosov flows from the perspective of the associated group actions on their orbit spaces and boundary at infinity. We extend the definition of Anosov-like action from [BFM25b] from the transitive to the general nontransitive context and show that one can recover many dynamical features such as the basic sets of a flow, and the Smale order on basic sets, from such general group actions.
Using these tools, we prove that a pseudo-Anosov flow on a compact $3$ manifold is determined by the associated action of the fundamental group on the boundary at infinity of its orbit space, answering an open question. We also give a number of other applications, including a proof that any topological pseudo-Anosov flow on an atoroidal 3-manifold is necessarily transitive, and that density of periodic orbits implies transitivity. These latter results fill some gaps in the literature, where proofs have only been written in the smooth Anosov case.
@@ 1. Introduction
This paper is motivated by the structure theory and classification of pseudo-Anosov flows in dimension 3. In [Mos92], Mosher introduced pseudo-Anosov flows as flows “locally modeled on the suspension of a pseudo-Anosov homeomorphism of a surface.” Such flows are characterized by the presence of a pair of transverse 2-dimensional invariant foliations, generalizing the weak-stable and unstable foliations of an Anosov flow, which Mosher suggested should be considered their defining feature. Since then, pseudo-Anosov flows have become increasingly recognized as important objects in hyperbolic dynamics and low-dimensional topology, and Mosher’s definition has been formalized in a variety of ways, resulting in definitions of “smooth pseudo-Anosov” or “topological (pseudo) Anosov” flow. However, there has not been widespread agreement on the definition, and the foundations remain incomplete. [^1] See [BM25, Remark 1.1.11.] for a discussion, and Example 1.2.15 of [BM25] for a cautionary example illustrating a critical gap in some existing definitions. Remark 2.1 below also gives further references.
@@
In this paper we develop a formal, axiomatic approach to the topological study of pseudo-Anosov flows, using the framework of Anosov-like actions of discrete groups on the plane. This perspective was introduced in [BFM25b] to give an invariant towards the classification of transitive pseudo-Anosov flows. The present paper develops further groundwork, applicable also to the nontransitive case. Using this we prove a number of structural results. For instance, we show that any (topological) pseudo-Anosov flow or topological Anosov flow on an atoroidal 3-manifold is necessarily transitive. This fills an apparent gap in the literature, since the existing proofs for smooth Anosov or pseudo-Anosov flows (due to Brunella [Bru93] and Mosher [Mos92]) use Smale’s spectral decomposition theorem, which is only fully proved in the smooth setting. We also characterize transitive versus nontransitive flows (see Theorem 1.5), and give many examples illustrating the nuances that appear in the presence of prongs in the pseudo-Anosov setting.
Finally, the main result of this article shows pseudo-Anosov flows are completely determined up to orbit equivalence (time-change and conjugacy) by an action of the fundamental group of the supporting manifold on $S^{1}$. This improves a theorem of Barbot, who showed such flows are determined by an action on $\mathbb{R}^{2}$, and fully settles what was an important question about the natural action on $S^{1}$ induced by pseudo-Anosov flow, which had been circulating since Fenley’s construction of the ideal $S^{1}$ boundary in 2005.
The orbit-space perspective
A pseudo-Anosov flow $\phi$ on a 3-manifold $M$ gives rise to an orbit space $\mathcal{O}_{\phi}$, a topological plane with two transverse (possibly singular) foliations which admits a natural action of $\pi_{1}(M)$ by homeomorphisms. Following work of Barbot [Bar95a], this action of a group of the plane (up to conjugacy) uniquely determines the flow up to orbit equivalence. Barbot’s theorem indicates that all features of a pseudo-Anosov flow and its orbit equivalence class should, in theory, correspond to features of the orbit space with the action of $\pi_{1}(M)$. Building such a “dictionary” between flows and their orbit spaces is now regarded as an essential tool in the subject. Early work of Barbot and Fenley used this idea towards a structure theory of pseudo-Anosov flows on graph manifolds [Bar96, BF13, BF15] among other results; more recently in [BFM25b, BFM25a] it was used to obtain a complete algebraic invariant of transitive pseudo-Anosov flows (on any given manifold $M$) via the dynamics of $\pi_{1}(M)$ on $\mathcal{O}_{\phi}$.
To develop a formal framework for pseudo-Anosov flows, we abstract the essential dynamical features of actions on orbit spaces, called an Anosov-like action. This allows one to use 2-dimensional topological dynamics to prove results about discrete groups acting on the plane, then translate this back to the 3-manifold setting.
Statement of results
A bifoliated plane is a topological plane $P$ with two transverse topological foliations $\mathcal{F}^{\pm}$, possibly with isolated prong singularities. The following definition captures the dynamics of the orbit-space actions of flows.
@@
We note that, for the results of this article, Axiom (A5) could be weakened slightly, requiring that we only forbid such quadrilaterals that do not contain any fixed points of elements of $G$ (see Remark 4.19). This weak version prevents examples of groups which split as free products (see Theorem 8.11), a behavior very different from the fundamental groups of manifolds supporting flows which are always aspherical. Proposition 2.2 below shows that the orbit space action from a pseudo-Anosov flow satisfies these axioms; however, these are not the only examples.
These axioms have surprisingly strong consequences, both for the topological structure of the bifoliated plane, and the global and local dynamics of the action of $G$. For example, we show (following Barbot) that a bifoliated plane admitting an Anosov-like action either has prongs or nonseparated leaves, is trivial, or has a special form called skew (Proposition 2.10). We also show a number of relationships between dynamics and group structure:
Theorem 1.2 (Point stabilizers).
@@
The proof does not require any 3-manifold topology (beyond showing that the orbit space of a flow is Hausdorff), nor require any smoothness of the flow, so applies also to topological (pseudo)-Anosov flows in the sense of [BM25, Def. 1.1.10] or [AT24, Def. 5.9]. We also show
@@
Again, this is well known for smooth Anosov flows, but to our knowledge has not been proved in the topological and pseudo-Anosov setting.
Remark 1.6.
As a consequence of the axioms, if a leaf $l$ has nontrivial stabilizer, then all nontrivial elements of the stabilizer have a common, unique fixed point in $l$. For singular leaves, this is a consequence of Axioms (A1) and (A3). In the nonsingular case, this follows from Axiom (A1) together with Solodov’s Theorem (see [BFM25b, Lemma 2.6]). Thus, Theorem 1.2 has an equivalent statement in terms of stabilizers of leaves.
An important tool in Theorem 1.3 and 1.5 as well as the general structure theory is the definition and description of Smale classes.
Smale classes and chains. Brunella [Bru93] showed that the basic sets of an Anosov flow are separated by a collection of tori transverse to the flow. In the nontransitive case any basic set contains “boundary” orbits, which are periodic orbits such that (at least) one of their rays are contained in the wandering set. A related theory is developed for Smale diffeomorphisms of surfaces in [BL98].
Here we generalize this theory to pseudo-Anosov flows, where the situation is significantly more subtle. One issue is that, in this setting, non-wandering components can be strictly finer than chain-recurrent components (see Proposition 8.8 for an example). This gives rise to unusual dynamical behavior not seen in the Anosov setting. To address this, we introduce a structural theory of Smale classes, which generalize the notion of a basic set for an Anosov flow. These inherit a natural partial order, which we denote by $\leq_{G}$ (see Definition 3.4). We show:
@@
This answers a problem from [BFM25b, Remark 2.4], see Corollary 1.11.
By [Bru93], when $M$ is orientable and compact, any two basic sets of an Anosov flow are separated by embedded tori in $M$ transverse to the flow. We describe the counterpart structure in the orbit space. These are called Smale chains. We show:
@@
See Theorem 4.23 for a more precise statement, in which we clarify what is meant by separating. Together with Theorem 5.5, this shows that, provided an Anosov-like action is cocompact, distinct Smale classes can be separated by $\mathbb{Z}^{2}$-invariant wandering chains.
Translating this to the language of pseudo-Anosov flows, we show that wandering chains separate distinct non-wandering components; whereas the (often coarser) chain-recurrent components are those separated by Lyapunov functions.
“Reduction of dimension”: classification by actions on $S^{1}$.
A bifoliated plane comes equipped with a natural compactification by a boundary at infinity (see [Bon24, Fen12, Mat82]) to which any foliation-preserving action by homeomorphisms extends. This compactification has become an important tool for the study of group actions on such planes and, as a special case, the study of pseudo-Anosov flows. It was shown in [BFM25b] that for transitive Anosov-like actions, the induced action at infinity uniquely determines the action on the interior. Here, using tools from [BBM25], we show this is true generally, with an independent proof.
@@
Let $G$ be a Smale-bounded Anosov-like action on a bifoliated plane $P$. The induced action on $\partial P$ uniquely determines the action of $G$ up to conjugacy.
The definition of Smale-bounded is given in Definition 6.1. For the purpose of this introduction, it is enough to know that every Anosov-like action with finitely many Smale classes is Smale-bounded (Proposition 6.3).
Consequently, pseudo-Anosov flows on compact 3-manifolds are determined up to orbit equivalence by the action of $\pi_{1}(M)$ on $\partial P$:
Theorem 1.10 (Circle actions classify flows).
The action of $\pi_{1}(M)$ on $S^{1}$, obtained from the compactification of the orbit space by a circle at infinity, (up to conjugacy) completely determines a pseudo-Anosov flow up to orbit equivalence.
The proof of this theorem has two ingredients: One is the work of the same authors done in [BBM25], which allows to uniquely recover the bifoliated plane (and, by uniqueness, also the group action) from relatively “sparse” data on the boundary (Corollary E of [BBM25], restated as Theorem 7.5 below). The second ingredient, obtained in this article, consists in proving that when the foliations are not $\mathbb{R}$-covered, the non-corner fixed leaves are dense in $P$. The endpoints of such leaves are sparse – it is a countable set – but sufficiently rich so that we can apply the result of [BBM25]. A non-corner fixed leaf is a leaf $l$ of $\mathcal{F}^{\pm}$, fixed by some nontrivial element $g\in G$, and such that the fixed point $x$ of $g$ on $l$ is not the corner of any lozenge (see Definition 2.5). Such non-corner fixed points and fixed leaves can be “seen” by the action of $G$ on the circle at infinity: an element $g$ fixes a non-corner leaf if and only if the subgroup generated by $g$ has exactly $4$ fixed points on $\partial P$ such that two are attracting and two are repelling (see Propositions 7.1 and 7.2). Thus, having two actions that are conjugated on the circle at infinity forces the non-corner fixed leaves for both actions to be “the same”, and the result of [BBM25] applies.
Thus, the technical work toward the proof of Theorem 1.9 is to show the density of non-corner fixed leaves. This is obtained in Corollary 6.11. In the case of topologically transitive Anosov-like actions, this is comparatively easy, and was obtained in [BFM25b]. In the nontransitive case however, it relies on the description of Smale classes and Smale chains that we obtain here.
@@
The axioms for transitive Anosov-like actions. Theorem 1.7 answers the question posed in [BFM25b, Remark 2.4], and in fact shows the axioms given there for transitive Anosov-like actions are redundant and can be simplified. To explain this, we recall that [BFM25b] gave a list of strictly stronger axioms than those in Definition 1.1 for groups acting on bifoliated planes, capturing the behavior of transitive pseudo-Anosov flows. In the transitive case, Axiom (A2) does not appear, and in its place we have the following two axioms.
@@
We point out that we also weakened slightly Axiom (A4) compared to [BFM25b]. We show in Proposition 2.12 that the weaker axiom of this article imply the stronger one of [BFM25b].
@@
Section 2 contains background on bifoliated planes and Anosov-like actions. We extend some results from [BFM25b] to the nontransitive case, as well as introduce definitions and a toolkit used later in this work.
@@
Section 5 gives the proof of Theorems 1.2 and 1.3 and improves Theorem 1.8 in the case of cocompact actions.
@@ 2. Preliminaries: Anosov-like actions and bifoliated planes
As mentioned in the introduction, our definition of Anosov-like actions on bifoliated planes is motivated by the dynamics of the action of the fundamental group of a 3-manifold on the orbit space of a pseudo-Anosov flow. If $\phi$ is a flow on a manifold $M$, the orbit space $\mathcal{O}_{\phi}$ of $\phi$ is the quotient of $\widetilde{M}$ by orbits of the lift of $\phi$ to $\widetilde{M}$. Fenley [Fen94] and Barbot [Bar95a] showed that, if $\phi$ is a (pseudo)-Anosov flow [^2] The case of pseudo-Anosov flow was treated by Fenley–Mosher in [FM01]. on a 3-manifold, then $\mathcal{O}_{\phi}$ is a topological plane. The weak stable and unstable foliations descend to transverse 1-dimensional singular foliations $\mathcal{F}^{\pm}$ on $\mathcal{O}_{\phi}$ preserved by the action of $\pi_{1}(M)$.
@@
There are two common notions of Anosov and pseudo-Anosov flows in the literature: smooth and topological (pseudo)-Anosov flows (see, e.g., [AT24], although even these definitions have some variation within them in different texts). These notions are only known to coincide, up to orbit equivalence, for transitive pseudo-Anosov flows ([Sha21, AT24]). The following proposition, and thus all the results of this article, applies even for the more general topological version of the definition (equivalent to the notion of an expansive flow – see [BM25, Section 1.1].)
@@ 2.1. Structure of the bifoliated planes
Following work of Barbot and Fenley, there is a well developed structure theory for the orbit spaces of Anosov flows. Several of these foundational results were established for bifoliated planes admitting (transitive) Anosov-like actions in [BFM25b]. Below, we recall some of this structure theory and establish necessary extensions to our setting.
@@
A ray of $\mathcal{F}^{\pm}(x)$ is a properly embedded copy of $0,\infty)$ in $\mathcal{F}^{\pm}(x)$ based at xx. [^3] Elsewhere these are also called half-leaves, we prefer the term ray because it is more accurate in the case of singular points.
@@
In other words, a closed lozenge is homeomorphic to a rectangle with two opposite corners removed, with the segments of $\mathcal{F}^{+}$-leaves (resp. $\mathcal{F}^{-}$) sent to horizontal (resp. vertical) segments. A basic and elementary fact is as follows:
Lemma 2.6 (Lemma 2.23 [BFM25b]).
If $x$ is fixed by some nontrivial $g\in G$ and some ray through $x$ makes a perfect fit with a leaf $l$, then this perfect fit is part of a lozenge with corner $x$, and opposite corner on $l$.
A point that is not the corner of any lozenge is called a non-corner point. If this point is fixed by some element $g\in G$, we call this a non-corner fixed point. These will play an important role later on. For now, we note for future reference (without proof) the following easy result:
Lemma 2.7 (Lemma 2.29 [BFM25b]).
Suppose $r^{+}$ and $r^{-}$ are two rays bounding a quadrant $Q$ of $x$. If each of $r^{+}$ and $r^{-}$ intersect one of a pair of leaves making a perfect fit, or if they each intersect leaves of a singular point, then there are no lozenges in $Q$ with $x$ as a corner.
The statement in [BFM25b] is for transitive Anosov-like actions, but the only axioms used are in our definition here and so the proof applies verbatim.
Lozenges may share corners, leading to what is called a chain of lozenges.
@@
A chain of lozenges is a union of closed lozenges that satisfies the following connectedness property: for any two lozenges $L,L^{\prime}$ in the chain, there exist lozenges $L_{0},\dots,L_{n}$ in the chain such that $L_{0}, L=L$$^{\prime}=L_{n}$, and, for all $i$, $L_{i}$ and $L_{i+1}$ share a corner (and may or may not share a side).
@@
The following global structure result is proved for transitive Anosov-like actions in [BFM25b, Theorem 2.16], following the outline of [Bar95a, Théorème 4.1]. It also holds for general Anosov-like actions, with minor modifications to the proof, which we include below.
@@
    $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ is trivial.
@@
    $(P,\mathcal{F}^{+},\mathcal{F}^{-})$ is homeomorphic to the diagonal strip bounded by y=x1y=x-1 and $y=x+1$ in $\mathbb{R}^{2}$ with the horizontal and vertical foliations. Such planes are called skewed.
@@
We describe the outline of the proof, for completeness, and describe the required modification from the proof given in [BFM25b, Theorem 2.16].
@@
Let $\mathcal{L}^{-}$ and $\mathcal{L}^{+}$ denote the leaf spaces of $\mathcal{F}^{-}$, $\mathcal{F}^{+}$ respectively. Suppose that $P$ has no singular points and $\mathcal{L}^{+}$ is Hausdorff. Since it is a simply connected 1-manifold it is thus homeomorphic to $\mathbb{R}$. We need to show in this case that $\mathcal{L}^{-}\cong\mathbb{R}$ and the plane is either skewed or trivial. Following the argument from [BFM25b, Theorem 2.16] verbatim, one shows that, either both leaf spaces are $\mathbb{R}$ and the plane is trivial, or each leaf $l\in\mathcal{L}^{-}$ meets a bounded subset of leaves of $\mathcal{L}^{+}$. Fix an identification of $\mathcal{L}^{+}$ with $\mathbb{R}$ and for $l\in L^{-}$, let $\alpha(l)$ and $\omega(l)$ denote the infimum, and supremum, respectively, of the sets of leaves of $\mathcal{F}^{+}$ intersected by $l$.
We claim any leaf $l$ of $\mathcal{F}^{-}$ separates $\alpha(l)$ and $\omega(l)$ in $P$. This claim implies that $\mathcal{L}^{-}$ is Hausdorff, hence homeomorphic to $\mathbb{R}$, and thus $P$ can be seen as a subset of the plane obtained as the product of the leaf spaces $\mathcal{L}^{+}\times\mathcal{L}^{-}\simeq\mathbb{R}^{2}$ bounded by the graphs of the function $\alpha$ and $\omega$, as in [Fen94, Theorem 3.4], this is easily seen to be homeomorphic to the skewed plane.